OpenStudy (anonymous):

The fraction 25/100 can be reduced to 5/20, a fraction having the same value, by dividing both the numbers 25 and 100 by

OpenStudy (anonymous):

well what can you divide into 25 that can also be divided into 100

OpenStudy (nurali):

The fraction 25/100 can be reduced to 5/20, a fraction having the same value, by dividing both the numbers 25 and 100 by: The answer is 5. You would divide both 25 and 100 to get 5 and 20 25/100=5/20=1/4